# Training the Untrainable: Introducing Inductive Bias via Representational Alignment

### v1

@@ 4 Experiments
To systematically evaluate our approach, we categorize our experiments into the following groups of untrainability that our proposed method improves upon:
  1. 1.
    Untrainable Architectures: Experiments where the target networks are difficult to train due to architectural limitations, irrespective of the task.
  2. 2.
    Untrainable Tasks: Experiments where certain tasks are inherently challenging for specific architectures, making them untrainable without additional supervision.
4.1 Tasks
Here, we describe the task settings. For an image-based task, we focus on image classification and use the ImageNet-1K dataset (Deng et al., 2009) for training and testing. We use the splits defined by the dataset. We report accuracy on the validation set for all experiments.
We next consider three sequence modeling tasks that allow for a broader range of architectural settings. We first consider a task called copy-paste (Graves, 2014). In this task, we generate a sequence of numbers in the range of $1-10$. The model is trained to recover the same sequence in the output. In our setting, we consider sequence lengths that range from $20$ to $40$ values total (internal sequence and padding). We generate a copy-paste dataset, sampling sequences containing numbers between $1$ and $10$. We generate a total of $100,000$ examples, training on $80,000$ examples, validating on $10,000$ examples, and testing on $10,000$ examples.
We also include the parity task, a binary classification task where a model is fed a bitstring and outputs $1$ when there is an even number of ones in the bitstring and 00 otherwise. We generate a series of bitstrings with sequence lengths that range from $2$ to $50$ (as in prior work) (Bhattamishra et al., 2020).
Finally, we consider a language modeling task using the WikiText-103 dataset (Merity et al., 2016) where models must predict the next token given some context. This uses the train, validation and testing splits defined by the WikiText dataset and for all experiments, we use a context length of $50$. We tokenize the text data using the GPT-2 (Radford et al., 2019) tokenizer.
4.2 Architectures
@@
Copy-Paste Transformer RNN
Parity RNN Transformer
Language Modeling Transformer RNN
@@
Table 1: Guide and target networks across tasks. Our network designs include several untrainable target networks and corresponding trainable guide networks. We focus on FCNs and deep convolutional networks for image classification and recurrent networks for sequence modeling.
@@
For all tasks, we describe our target untrainable architectures for each task separately as well as the guide networks that are employed to make the untrainable network trainable. We give an overview in table 1.
Image Classification: We use three target networks: Deep FCN, Wide FCN, and Deep ConvNet. Deep FCN is a fully-connected network with 50 blocks consisting of feedforward layers followed by non-linearities. This network is an untrainable architecture, lacking inductive biases to prevent overfitting and having vanishing gradients. Wide FCN is a fully connected network with 3 blocks with feedforward layers that have 8192 units. This is categorized as an untrainable task due to a saturation in the training performance. Deep ConvNet is the same architecture as ResNet-50 (He et al., 2016a), but without residual connections. This is categorized as an untrainable architecture due to the vanishing gradient problem. We use two guide networks: ResNet-18 and ResNet-50. ResNet-18/50 is a deep convolutional network with 18/50 convolutional blocks and residual connections. We refer to He et al. (2016a) for ResNet 18/50 design. We supervise the Deep FCN and Shallow FCN with ResNet-18 and supervise the Deep ConvNet with ResNet-50.
@@
Sequence Modeling: For our copy-paste task, we use a vanilla, 4-layer RNN and language modeling tasks, we use a vanilla 5-layer RNN as our target networks. In copy-paste, architectural and algorithmic limitations make RNNs an untrainable architecture. For language modeling, vanishing gradients and limited context incorporation make RNNs an untrainable architecture as the training loss saturates. For the parity task, we use a 1-layer transformer encoder architecture, similar to prior work (Bhattamishra et al., 2020; Hahn & Rofin, 2024). For the copy-paste task, we train a guide network, 4-layer transformer decoder model which achieves 96.90% accuracy. Similarly, for language modeling, we train a 5-layer transformer decoder guide network with a context window of 256. Our final perplexity is 34.15. For the parity task, we train a 1-layer vanilla RNN as a guide network which achieves the same 100% accuracy reported by Bhattamishra et al. (2020).
4.3 Training
For each setting, we train the base target network and perform an experiment where both a trained and untrained guide network supervises the base target network. All networks are trained with either cross-entropy loss or binary cross-entropy loss, although other loss functions could be applied as well.
@@ 4.3 Training
When training networks for object detection using ImageNet-1K, we use the Adam (Kingma, 2014) optimizer. For all sequence modeling tasks, i.e. copy-paste, parity, and the language modeling task we use AdamW (Loshchilov, 2017). For language modeling, we also incorporate gradient clipping due to unstable training with long sequences.
@@
To ensure consistency of comparisons across learning curves, we use a consistent batch size of 256. In general, representational similarity metrics are affected by the number of samples in the calculation. The more samples, the better the metric approximates representational distance. We use 256 as a proxy, dependent on GPU memory, although more memory would allow for bigger batch sizes with potentially better results. Due to the large number of training settings, we employ several different learning rates. We tune the learning rate carefully for baseline training to ensure maximal performance of base training. We sweep the parameter across 5 different values and choose the results with the lowest validation loss. This ensures we are choosing the training with the best performance.
@@
After choosing the optimal learning rate, we then train all networks and settings for 100 epochs with 5 random seeds to compute error bars for our training loss, validation loss, and testing accuracy. Our error bars are associated with the standard error across each step across all seeds. We choose the seed-based average test accuracy associated with the epoch with the lowest seed-based average validation loss.
@@ 5 Results
Here, we report the findings from our use of guidance to train our set of untrainable networks. Results are shown in the following order: image classification task; sequence modeling tasks; language modeling tasks.
Experiment ImageNet Top-5 Validation Accuracy ()($\uparrow$)( ↑ )
@@
Experiment Copy-Paste Accuracy ()($\uparrow$)( ↑ ) Parity Accuracy ()($\uparrow$)( ↑ ) Language Modeling Perplexity ()($\downarrow$)( ↓ )
RNN 14.35 $\pm$ 0.01 100 69.19 $\pm$ 1.89
Untrained RNN 2.32 $\pm$ 0.41
Transformer 96.98 71.98 $\pm$ 3.16 34.15
Untrained Transformer 1.04 $\pm$ 0.81 51948.8 $\pm$ 90.44
RNN $\rightarrow$ Transformer 78.49 $\pm$ 2.16
Untrained RNN $\rightarrow$ Transformer 70.38 $\pm$ 4.17
Transformer $\rightarrow$ RNN 23.27 $\pm$ 1.02 40.01 $\pm$ 1.54
Untrained Transformer $\rightarrow$ RNN 42.56 $\pm$ 1.51 59.61 $\pm$ 2.33
Table 3: Guidance improves performance for sequence modeling. RNN performance improves dramatically when aligning with the representations of a Transformer for copy and paste, as well as for language modeling. RNNs close most of the gap to Transformers for language modeling. Transformers in turn, improve parity performance when aligning with an RNN. Guidance is able to transfer priors between networks.
5.1 Image Classification
Refer to caption Deep FCNRefer to caption Wide FCNRefer to caption Deep ConvNetRefer to caption Copy-Paste, RNN TargetRefer to caption Parity, Transformer TargetRefer to caption Language Modeling, RNN Target
Figure 2: Training and validation under guidance for all experiments reported in table 1. For every result in Table 2 and Table 3 we show the training and validation loss with error bars across multiple runs although these are often too small to see. Note that often the best results occur with the untrained guide.
Across all our networks, we observe significant improvement from using a guide network to provide guidance; see table 2. We see significant accuracy gains of 5-10% on validation performance. We also observe significantly better loss curves from a better fit with the training loss and reduced overfitting with the validation loss. Most interestingly, we highlight that using a randomly initialized guide network can perform better than using a trained guide network. For example, the Deep FCN results in the top left of fig. 2 are significantly better with a randomly initialized ResNet-18 as the guide network instead of a trained ResNet-18. This trend also occurs with Wide FCN.
We note that this trend is not entirely consistent as indicated by the Deep ConvNet. One potential explanation is that the architectural prior is the same for both the Deep ConvNet and ResNet-50. This explanation provides an additional interpretation for the role of residual connections and their influence on the representation space. This indicates that residual connections must be trained to have an influence on the representation space, and aligns with prior studies of residual connections (Jastrzebski et al., 2017; He et al., 2016b).
5.2 Sequence Modeling
On the copy-paste task, see fig. 2 and table 3, we see significant improvement when using a transformer as our guide network over an RNN target network. Prior work has observed the difficulty that RNNs have with copy-paste and usually attributes this to a fundamental algorithmic limit on the capability of RNNs to memorize input sequences. Our results show a potential optimization scheme for RNNs that is applicable for sequence memorization. Similar to prior results with fully connected networks, this result improves when using a randomly initialized network. We explore this more thoroughly in further analyses but this is likely because optimization with randomly initialized networks is easier due to the degrees of freedom in CKA.
@@ 5.2 Sequence Modeling
On the parity task, we similarly see improvement when using an RNN as our guide network over a transformer model as our target network. This is a complementary result to copy-paste and language modeling where the guide network was a transformer and our target network was a RNN. This improves over results from several prior papers that have pointed out fundamental limitations of transformers to perform certain formal language tasks.
@@
Unlike the prior tasks, the performance improves when using a trained RNN as the guide network. This could be due to the wide gap in performance between an untrained RNN and trained RNN on parity. Trained RNNs achieve 100% on the task and could have meaningful positional information to complete the parity task. This information is likely crucial to the transformer which has limited sequence pooling capacity and fewer degrees of freedom unlike an RNN.
@@
On language modeling, similar to copy-paste, we see significant improvements with the RNN target network that is guided by a transformer guide network. While performance generally saturates for the 5-layer RNN, guide network guidance continuously improves the RNN performance by close to 30 points for test perplexity for both trained guide networks and 10 points for randomly initialized guide networks. This implies that information from the transformer can be transferred to the RNN. We see a similar trend with a randomly initialized transformer as the guide network implying that architectural priors in the transformer are driving much of the improvement in guided network performance.
5.3 Representational Similarity Loss
Refer to captionRefer to caption
Figure 3: Representational dissimilarity changes across training. We view the representational dissimilarity, measured as the complement of CKA, across training between our target and guide networks. We find that across many settings, target networks reduce dissimilarity more quickly with untrained guide networks, which can explain improved results with untrained guide networks. With language modeling, we find that the RNN reduces dissimilarity with an untrained guide network more quickly despite having a worse performance.
We can view the representational alignment between the guide and target networks during training. This allows us to better understand how this representational alignment influences network performance. We show results in fig. 3.
We notice that across most tasks, reducing representational dissimilarity is easier with activations from randomly initialized networks rather than trained networks. This provides additional evidence of representational alignment for inductive bias transfer. We notice that for certain cases, such as Parity, the randomly initialized guide network has higher representational dissimilarity loss than the trained guide network. This is matched with the Parity result in table 3 and fig. 2.
However, we can also observe more inconsistent results with the Deep ConvNet where the untrained guide network has lower representational dissimilarity loss than the trained guide network, even at the end of training. One possible explanation here is that the inductive bias was more similar for Deep ConvNet and ResNet-50. This means that alignment with an untrained model is uninformative as the only information in an untrained guide network is its own inductive bias which Deep ConvNet matches more closely. So, it is likely that trained features are improving guidance with Deep ConvNet.This result has interesting implications for understanding the role of residual connections. The untrained residual connections have little influence on the inductive biases of the network or the overall representation space. This demonstrates the strength of our method as a way to interpret neural network design choices and how they influence representation and functional aspects of a network. Similarly, for the RNN trained for language modeling, we hypothesize that trained features are more useful for improving RNN performance for a similar reason.
These results are also potentially indicative of architectural properties of RNNs and FCNs which match randomly initialized networks more quickly. For instance, one potential explanation is that RNNs have more degrees of freedom (Bhattamishra et al., 2020) and therefore, only need inductive guidance rather than trained features. Transformers may require learned features indicating that the bottleneck for transformers on the parity is not algorithmic but feature-based. Future work can use these results as a starting point for the specification of more informative priors in model design (i.e. architectural priors) and training procedures alike.
5.4 Error Consistency
@@
Figure 4: Guidance aligns error consistency. The relationship between the guide networks is mirrored in that of the guided networks even when the target is entirely unlike the guides initially. This is additional evidence that guidance doesn’t just improve performance arbitrarily, the target becomes more like the guide.
[…]
$\kappa=\frac{c_{\text{obs}}-c_{\text{exp}}}{1-c_{\text{exp}}}$ (7)
$\kappa$ ranges from $-1$ to $1$, where $1$ is perfect agreement, $-1$ is perfect disagreement and 00 is change agreement. When $\kappa>0$, this implies that models make consistent error patterns, $\kappa<0$ implies that models make inverse error patterns, and $\kappa\approx 0$ implies independent error patterns.
@@
See fig. 4. The error consistency between the initial FCNs is entirely unlike the ResNet-18 or ViT-B. Guidance creates two FCNs which have the same relationship to one another. The ResNet-18-guided FCN and ViT-B-guided FCN have the same error consistency with respect to one another as ResNet-18 and ViT-B do. In other words, it’s not just that the FCN gets generically better; rather, it’s that the FCN adopts a prior from the original architecture.
5.5 Network Initialization
Figure 5: Disconnecting the guide early can have long-term effects. We remove the ResNet guide from the loss function of a Deep FCN after an arbitrary 150 steps. Surprisingly, this has little effect. The Deep FCN still does not diverge and the training and validation loss still mirror one another. Very likely, an initialization regime exists for Deep FCNs that prevents overfitting. We now have the tools to investigate this question.
@@ 5.5 Network Initialization
Is guidance needed throughout training, or is the effect of the guide to move the target into a regime where the two are aligned and the target can be optimized further without reference to the guide? The answer to this question can shed light on whether the guide is telling us that better initializations are likely to exist for the target. To answer this question, we disconnect the guide from the target after an arbitrary and nominal number of training steps, 150. We did not investigate what the minimal number of steps needed was.
For FCNs, even this early disconnect avoids overfitting; see fig. 5. Furthermore, while preventing overfitting, we have lower training loss from guidance, indicating a better fit. This implies that there likely exists a better initialization for FCNs. In future work, we intend to investigate what the untrained guide is doing to the FCN weights to prevent validation loss from increasing dramatically.

### v2

@@ 4 Experiments
To systematically evaluate our approach, we incorporate two settings. (1) Untrainable Architectures: Experiments where the target networks are difficult to train due to architectural limitations, irrespective of the task. For example, memory incorporation in RNNs or overfitting in deep FCNs. (2) Untrainable Tasks: Experiments where certain tasks are inherently challenging for specific architectures, making them untrainable without additional supervision. For example, sequence classification with transformers.
Tasks: We describe the task settings. We consider three sequence modeling tasks to allow for a broader range of architectural settings. We first consider a task called copy-paste [27]. In this task, we generate a sequence of numbers in the range of 1101-10. The model is trained to recover the same sequence in the output. In our setting, we consider sequence lengths that range from $20$ to $40$ values total (internal sequence and padding). We generate a copy-paste dataset, sampling sequences containing numbers between $1$ and $10$. We generate a total of $100,000$ examples, training on $80,000$ examples, validating on $10,000$ examples, and testing on $10,000$ examples.
We also include the parity task, a binary classification task where a model is fed a bitstring and outputs $1$ when there is an even number of ones in the bitstring and 0 otherwise. We generate a series of bitstrings with sequence lengths that range from $2$ to $50$ as done in prior work [8].
Finally, we consider a language modeling task using the WikiText-103 dataset [54] where models must predict the next token given some context. This uses the train, validation and testing splits defined by the WikiText dataset and for all experiments, we use a context length of $50$. We tokenize the text data using the GPT-2 [61] tokenizer.
For an image-based task, we focus on image classification and use the ImageNet-1K dataset [19] for training and testing. We use the splits defined by the dataset. We report accuracy on the validation set for all experiments.
@@
Copy-Paste Transformer RNN
Parity RNN Transformer
Language Modeling (Small and Large) Transformer RNN
@@
Table 1: Guide and target networks across tasks. Our network designs include several untrainable target networks and corresponding trainable guide networks.
@@
Architectures: For all tasks, we describe our target untrainable architectures for each task separately as well as the guide networks that are employed to make the untrainable network trainable. We give an overview in table˜1. We provide further details in appendix˜C.
@@
Sequence Modeling: For our copy-paste task, we use a vanilla, 4-layer RNN as our target network. In copy-paste, architectural and algorithmic limitations make RNNs an untrainable architecture. For our language modeling task, we include two settings with a small (4 layer) and large (6 layer, larger hidden dimension) RNN. In this setting, vanishing gradients and limited context incorporation make RNNs an untrainable architecture as the training loss saturates. For the parity task, we use a 1-layer transformer encoder architecture, similar to prior work [8, 28]. For the copy-paste task, we train a guide network, 4-layer transformer decoder model which achieves 96.90% accuracy. Similarly, for language modeling, we train a 4-layer transformer decoder guide network with a context window of 256. Our final perplexity is 34.15 for the small language modeling setting and 33.10 for the large language modeling setting. For the parity task, we train a 1-layer vanilla RNN as a guide network which achieves 100% accuracy as reported by [8].
Image Classification: We use three target networks: Deep FCN, Wide FCN, and Deep ConvNet. Deep FCN is a fully-connected network with 50 blocks consisting of feedforward layers followed by non-linearities. This network is an untrainable architecture, lacking inductive biases to prevent overfitting and having vanishing gradients. Wide FCN is a fully connected network with 3 blocks with feedforward layers that have 8192 units. This is categorized as an untrainable task due to a saturation in the training performance. Deep ConvNet is the same architecture as ResNet-50 [31], but without residual connections. This is categorized as an untrainable architecture due to the vanishing gradient problem. We use two guide networks: ResNet-18 and ResNet-50. ResNet-18/50 is a deep convolutional network with 18/50 convolutional blocks and residual connections. We refer to He et al. [31]. We supervise the Deep FCN and Wide FCN with ResNet-18 and supervise the Deep ConvNet with ResNet-50.
@@
Training: For each setting, we train the base target network and perform an experiment where both a trained and untrained guide network supervises the base target network. All networks are trained with cross-entropy loss, without loss of generality. For all sequence modeling tasks, i.e. copy-paste, parity, and language modeling we use AdamW [52]. For language modeling, we also incorporate gradient clipping due to unstable training with long sequences. When training networks for image classification using ImageNet-1K, we use the Adam [43] optimizer.
@@
To ensure consistency of comparisons across learning curves, we use a consistent batch size of 256. Representational similarity metrics are affected by the number of samples in the calculation, where more samples allows for the metric to approximate representational distance better. We use 256 as a proxy, dependent on GPU memory, although more memory would allow for bigger batch sizes with potentially better results. Due to the large number of training settings, we employ several different learning rates. We tune the learning rate carefully for baseline training to ensure maximal performance. We sweep the parameter across 5 different values and choose the results with the lowest validation loss. This ensures we are choosing the training with the best performance.
@@
After choosing the optimal learning rate, we then train all networks and settings for 100 epochs with 5 random seeds to compute error bars. Our error bars are associated with the standard error across each step across all seeds. We choose the seed-based average test accuracy associated with the epoch with the lowest seed-based average validation loss.
@@ 5 Results
Experiment
Copy-Paste
Accuracy $(\uparrow)$
Parity
Accuracy $(\uparrow)$
Language Modeling
(Small) Perplexity()($\downarrow$)
Language Modeling
(Large) Perplexity $(\downarrow)$
RNN 14.35 $\pm$ 0.01 100 69.19 $\pm$ 1.89 89.13 $\pm$ 2.00
Untrained RNN 2.32 $\pm$ 0.41
Transformer 96.98 71.98 $\pm$ 3.16 34.15 33.10
Untrained Transformer 1.04 $\pm$ 0.81 5.19e5 $\pm$ 90.44 5.19e5 $\pm$ 90.44
RNN $\rightarrow$ Transformer 78.49 $\pm$ 2.16
Untrained RNN $\rightarrow$ Transformer 70.38 $\pm$ 4.17
Transformer $\rightarrow$ RNN 23.27 $\pm$ 1.02 40.01 $\pm$ 1.54 36.91 $\pm$ 1.04
Untrained Transformer $\rightarrow$ RNN 42.56 $\pm$ 1.51 59.61 $\pm$ 2.33 47.17 $\pm$ 2.50
Table 2: Guidance improves performance for sequence modeling. RNN performance improves dramatically when aligning with the representations of a Transformer for copy and paste, as well as for language modeling with small and large RNN architectures. RNNs close most of the gap to Transformers for language modeling and are likely competitive with further scale. Transformers in turn, improve parity performance when aligning with an RNN. Guidance is able to transfer priors between networks.
Experiment ImageNet Top-5 Validation Accuracy $(\uparrow)$
@@
Sequence Modeling: On the copy-paste task, guiding a 4-layer RNN with a Transformer improves copy-paste accuracy by over 25%. See fig.˜2 and table˜2 Previous studies blamed RNN failures on vanishing gradients and memorization limits. Our results show a potential optimization scheme for RNNs that is applicable for sequence memorization. Remarkably, a random (untrained) Transformer guide outperforms a trained one, suggesting pure architectural bias drives gains. We believe this is because optimization with randomly initialized networks is easier due to the degrees of freedom in CKA. See appendix˜E. We plan to explore this more thoroughly in further analyses. These gains persist under our layerwise ablations and metrics; see appendix˜K and appendix˜G.
@@
On the parity task, a 1-layer Transformer guided by an RNN improves its test accuracy by 7%. This is a complementary result to copy-paste and language modeling where the guide network was a transformer and our target network was a RNN. This improves over results from several prior papers [8] that have pointed out fundamental limitations of transformers to perform formal language tasks.
@@
Unlike copy-paste, the performance improves when using a trained RNN as the guide network. This could be due to the wide gap in performance between an untrained RNN and trained RNN on parity. Parity uniquely benefits from learned positional encodings in the trained RNN, which the Transformer lacks. This information is likely crucial to the transformer, which has limited sequence pooling capacity and fewer degrees of freedom.
@@
On language modeling, similar to copy-paste, guided RNNs halve the perplexity from ~70 points to 35 points on WikiText-103, closing in on Transformer baselines. While performance generally saturates for the 4-layer RNN, guidance continuously improves the RNN performance by over 30 points for text perplexity for both trained and randomly initialized guide networks. Scaling up to a 6-layer RNN further cuts perplexity by 10 points, indicating guidance scales with model size. This implies that information from the transformer can be transferred to the RNN. We also believe that this has exciting implications for scaling laws with RNNs. We see a similar trend with a randomly initialized transformer as the guide network, implying that architectural priors in the transformer are driving improvement in guided network performance.
Figure 2: Training and validation under guidance for all experiments reported in table˜1. For every result in Table˜3 and Table˜2, we show the training and validation loss with error bars across multiple runs, although these are often too small to see. Note that often the best results occur with the untrained guide.
Image Classification: Guidance boosts validation accuracy by 5–10% across our Deep FCN, Wide FCN, and Deep ConvNet; see table˜3. We also observe significantly better loss curves from a better fit with the training loss and reduced overfitting with the validation loss. An untrained ResNet-18 guide outperforms its trained counterpart on Deep FCN, underscoring pure architectural priors like with copy-paste. For example, the Deep FCN results in the top left of fig.˜2 are significantly better with a randomly initialized ResNet-18 as the guide network instead of a trained ResNet-18. This trend also occurs with Wide FCN. We show results with other neural distance functions in appendix˜J.
The Deep ConvNet (no skips) gains only from a trained ResNet-50 guide—implying residual connections require learned weights to shape representation. This explanation provides an additional interpretation for the role of residual connections and their influence on the representation space. This indicates that residual connections must be trained to have an influence on the representation space. This aligns with prior studies of residual connections [39, 32].
@@
Figure 3: Guidance aligns error consistency. The relationship between the guide networks is mirrored in that of the guided networks, even when the target is entirely unlike the guides initially. This is additional evidence that guidance doesn’t just improve performance arbitrarily; the target becomes more like the guide.
@@
Error Consistency: Guided FCNs mirror the ResNet–ViT error overlap, proving they inherit the guide’s decision patterns. Using Deep FCN as our target model, we guide it with a ResNet-18 or a ViT-B-16 [21]. We then measure the error consistency [23] between all of the networks; see fig.˜3. The error consistency between the initial FCNs is entirely unlike the ResNet-18 or ViT-B. Guidance creates two FCNs which have the same relationship to one another. The ResNet-18-guided FCN and ViT-B-guided FCN have the same error consistency with respect to one another as ResNet-18 and ViT-B do. It’s not just that the FCN gets generically better; it adopts a prior from the original architecture. We provide further details of the error consistency metric in appendix˜F.
Figure 4: Initializing fully connected networks with guidance can overcome overfitting. First, we align a Deep FCN to a random ResNet-18 on noise for 300 steps, then train normally. This two-stage scheme mirrors full guidance, and leads to a similar performance improvement. This suggests that FCNs have guidance-inspired initializations that avoid overfitting.
@@
Network Initialization: Is guidance needed throughout training, or is the effect of the guide to move the target into a regime where the two are aligned and the target can be optimized further without reference to the guide? The answer to this question can shed light on whether the guide is telling us that better initializations are likely to exist for the target. To answer this question, we minimize the representational dissimilarity between our target and guide network for a nominal number of training steps, 300. Then we apply task training on the resulting target network with no guidance. Pre-aligning FCN layers to a random guide for 300 steps stops overfitting entirely—no ongoing guidance needed; see fig.˜4. Furthermore, while preventing overfitting, we have lower training loss from guidance, indicating a better fit. This implies that there exists a better initialization for FCNs.
Figure 5: Guidance outperforms distillation: We include a comparison between guidance and distillation for all settings with trained and untrained guide networks/teacher networks. We find that guidance outperforms distillation in all settings, highlighting that, unlike guidance, distillation fails in settings with an untrained teacher.
Guidance vs Distillation: We include a comparison between guidance and knowledge distillation [33] in fig.˜5. We find that guidance improves significantly over distillation, particularly when the teacher network is untrained. This is significant, demonstrating that guidance can exploit untrained networks for transferring inductive biases. This indicates that matching internal representations provides a much stronger signal over just matching output behavior. We discuss further in appendix˜H.